OpenStudy (anonymous):

5) What is the slop of the line defined by the equation 8x+2y=5?

OpenStudy (mosaic):

Rewrite the equation in the form y = mx + b. 'm' will be the slope.

OpenStudy (anonymous):

what about the 2?

OpenStudy (mosaic):

8x+2y=5 Solve for y by isolating y. Subtract 8x from both sides: 2y = -8x + 5 divide by 2: y = -4x + 5/2 Compare it to y = mx + b. slope m = -4. y-intercept b = 5/2.